OpenStudy (anonymous):

The volume of a cylinder is 24 cubic feet. What is the volume of a cone that fits exactly inside the cylinder?

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

that means cone needs to have the same height and radius as cylinder

ganeshie8 (ganeshie8):

volume of cylinder = pir^2h volume of cone with same dimensions = 1/3 pir^2h = 1/3 * volume of cylinder

OpenStudy (anonymous):

so 1/3 (24)? that would give me 8

ganeshie8 (ganeshie8):

Yup !!

OpenStudy (anonymous):

Thank you so much..i understand this now.

ganeshie8 (ganeshie8):

np :) good job !

OpenStudy (anonymous):

i have one more question is you have the time.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

shoot...

OpenStudy (anonymous):

The volume of a cube is 36 cubic feet. What is the volume of a pyramid that fits exactly inside the cube? Would i do the same as the last question?

ganeshie8 (ganeshie8):

yup ! volume of prism = lwh volume of pyramid wid same base and height = 1/3*lwh

ganeshie8 (ganeshie8):

so just take 1/3 of the given volume

OpenStudy (anonymous):

1/3(36)? 10.8?

ganeshie8 (ganeshie8):

1/3(36) = 12

ganeshie8 (ganeshie8):

how did u get 10.8 ?

OpenStudy (anonymous):

thats right....3 time 12 is 36....haha

ganeshie8 (ganeshie8):

:)

OpenStudy (anonymous):

calculator malfunction

ganeshie8 (ganeshie8):

burn it !